OpenStudy (anonymous):

What are the coordinates of the point on the graph of f(x)=(x+1)(x+2) at which the tangent is parallel to the line with equation 3x-y-1=0?

OpenStudy (tkhunny):

Have you considered the first derivative?

OpenStudy (amorfide):

since it mentions paralell the gradient of f(x) will be equal to the gradient of 3x-y-1=0 (re arrange this into y=mx+b format) so you will need to differentiate f(x) then make it equal to the gradient of the linear equation then solve for x then you can work out y and you have your coordinates

OpenStudy (anonymous):

yes I first got f(x)=x^2+3x+2 and then the derivative f'(x)=2x+3

OpenStudy (anonymous):

ok I will try that

OpenStudy (anonymous):

Ok so I had the derivative f'(x)=2x+3 and then I had that equal to the derivative of the second equation which is y'=3. So 2x+3=3. Then i solved for x and for x=0. would that be correct so far?

OpenStudy (tkhunny):

f(x) = (x+1)(x+2) = x^2 + 3x + 2 f'(x) = 2x + 3 = Gradient of f(x) for any given value of x 3x - y - 1= 0 ==> y = 3x - 1 Gradient of Line (Slope) = 3 Solve: f(x) gradient = Line Gradient Solve: 2x + 3 = 3 ==> x = 0 This is the desired x-coordinate. f(0) = 2 This is the desire y-coordinate. Answer: (0,2) Try to be WAY more organized and WAY less haphazard.